# Deep clustering using adversarial net based clustering loss

### v2

@@ Abstract
Deep clustering is a recent deep learning technique which combines deep learning with traditional unsupervised clustering. At the heart of deep clustering is a loss function which penalizes samples for being an outlier from their ground truth cluster centers in the latent space. The probabilistic variant of deep clustering reformulates the loss using $KL$ divergence. Often, the main constraint of deep clustering is the necessity of a closed form loss function to make backpropagation tractable. Inspired by deep clustering and adversarial net, we reformulate deep clustering as an adversarial net over traditional closed form $KL$ divergence. Training deep clustering becomes a task of minimizing the encoder and maximizing the discriminator. At optimality, this method theoretically approaches the $JS$ divergence between the distribution assumption of the encoder and the discriminator. We demonstrated the performance of our proposed method on several well cited datasets such as SVHN, USPS, MNIST and CIFAR10, achieving on-par or better performance with some of the state-of-the-art deep clustering methods.
I Introduction
@@
Deep neural network such as the autoencoder is applied to many signal processing domains such as speech and audio processing [1, 2], image clustering [3, 4] and medical data processing [5, 6]. While the latent space of autoencoder is well suited for dimension reduction through reconstruction loss [7], the latter is not optimized for clustering/ classification since class labels cannot be used in the reconstruction loss [8]. A recent autoencoder technique known as the Euclidean distance based clustering (ABC) [7], utilize class label information in neural network by introducing a loss function known as the deep clustering loss. The goal is to minimizes the Euclidean distance in the latent space between samples and partitioning learnt by clustering algorithms e.g. Kmeans or Gaussian mixture model (GMM). In other words, the latent space of an autoencoder learnt using reconstruction loss will look different when we further perform K-means on the latent space. A recent probabilistic approach to ABC uses KL divergence (KLD) [4, 9] and assumes that both the variational autoencoder (VAE) [10] latent space and the unsupervised clustering approach are Gaussian distributed. We can compare the similarities and differences between deep clustering and VAE to better understand the former. The neural network of VAE and deep clustering are optimized by backpropagating samples in the latent space to update the encoder weights. Also, both deep clustering and VAE share a similar goal of modeling each input image as a sample draw of the Gaussian distribution in the latent space. However, their Gaussian distributions are different. VAE enforce samples in the latent space to be closely distributed to a Gaussian distribution with continuous mean $\mu$ and variance $\sigma$ i.e. $z\sim\mu+\sigma\cdot\mathcal{N}(0,1)$. Whereas the latent space of deep clustering approaches a Gaussian mixture model (GMM) and each Gaussian has a set of fixed parameters i.e. $z\mathcal{\sim N}(\eta_{k},\tau_{k})$, $\forall k\in K$. The goal of VAE is to enforce all class samples in the latent space to be distributed and centered around a zero mean and unit variance Gaussian, $\mathcal{N}(0,1)$. The goal of deep clustering is to enforce class samples in the latent space to be closely distributed to their respective nearest Gaussian clusters i.e. $N(\eta^{*},\tau^{*})$ and nearest cluster is denoted by $\eta^{*}=\arg\max_{k}\eta_{k}$ and vice versa for $\tau^{*}$. Thus, only deep clustering is associated with clustering, making it more suitable for unsupervised classification.
Deep clustering is not without problem. Mainly, when backpropagating from a neural network, we require a closed form loss function for tractability [11]. When we relax from the constrain of assuming a KLD between two Gaussians for deep clustering, closed form loss function is seldom available for backpropagating.
Two main contributions in this work are as follows:
i) Probabilistic approach to deep clustering: We propose a probabilistic approach to relate the Euclidean based deep clustering in [7] to a JS divergence (JSD) based deep clustering using adversarial net. We also show in the Appendix that the evidence lower bound based deep clustering in VaDE [9] can be reduced to the same objective as the KLD based deep clustering in [4].
ii) Deep clustering using adversarial net: Most works often associate adversarial net with generative adversarial network (GAN), However, is it too restrictive to assume adversarial net can only perform GAN. The applications of adversarial net is in fact versatile as shown in this paper. Instead, we should generalize adversarial net as an approximator to the JSD, where the discriminator can be performing a specific task different from the discriminators in other adversarial net. We distinguish the proposed work from other “adversarial net based X” in Table I where X can be any discriminator task. We name our approach as “Deep clustering using adversarial net” (DCAN). In GAN, X refers to the binary threshold on whether the generator can produce images that are identical to the grouth truth. On the other hand, the discriminator of DCAN performs a decision on whether the encoder output approaches a GMM/Kmeans partitioning. More specifically in DCAN, the discriminator penalizes the encoder for outputing samples that fall outside their respective Kmeans/GMM assigned clusters in the latent space. Other instances of “adversarial net based X” includes AAE, DASC and etc. In AAE, the discriminator performs a check on whether the encoder can perform VAE. In DASC, the discriminator checks whether the encoder can perform subspace clustering. More details are given in Section 1.1 and Table I.
Section 2 establishes some background behind deep clustering and adversarial net. We introduce deep clustering using a Euclidean distance based loss known as ABC. Next, we extend deep clustering using the probabilistic variant of ABC based on a loss that assumes a “KLD between two Gaussians”. Lastly, we discuss the underlying link between adversarial net and the “JSD between two distributions” and how they are related to deep clustering. In Section 3, we proposed DCAN, deep clustering using adversarial net based loss. Finally, in Section 4, we evaluate DCAN on some well-cited image classification datasets such as USPS, MNIST and CIFAR10. Experimental results shows that DCAN is able to achieve on-par or better performance with several state-of-the-art deep clustering methods.
I-A Related works on adversarial net based X
The optimization of adversarial net [11] in general centers around minimizing the encoder/generator and maximizing the discriminator. However, as mentioned “adversarial net based X” are versatile and we highlight their similarities and differences as follows: In GAN, the generator is responsible for generating an image when given a sample from the latent space. The discriminator of GAN checks whether the generated image belongs to the distribution of the original dataset. Similarly, in DCAN we minimize the encoder and maximize the discriminator but for different purpose. In DCAN, the encoder receives an image and outputs a sample in the latent space. GAN trains the generator to be sensitive to the variance of the image class i.e. small displacement of the sample in the latent space can result in huge image variance. DCAN trains the encoder to be insensitive to the class variance i.e. different images from the same class populate tightly close to the cluster center in the latent space. Both DCAN and AAE [12] utilize the adversarial net in totally different ways despite having similar architecture (encoder acting as generator). The goal of AAE is identical to VAE i.e. model a sample in the latent space as a sum and product of $\mathcal{N}(0,1)$. On the other hand, the goal of DCAN is not identical to VAE. DCAN models a sample in the latent space, as the sample draw of a Kmeans, specifically the sample draw of a $1-of-K$ cluster. The goal of DASC [13] is subspace clustering, and it is different from deep clustering. Like principal component analysis, DASC perform unsupervised learning by decomposing the dataset into different individual eigenvectors, each eigenvector as orthogonal to each other as possible. Samples from each classes should ideally reside on their respective $K^{th}$ subspaces, just as each sample should ideally fall within their respective $K^{th}$ clusters in deep clustering. In DASC, the latent space of DASC is trained to behave like subspace clustering. Thus the adversarial net of all four methods, GAN (models the dataset), AAE (models VAE loss), DASC (models subspace clustering) and DCAN (models deep clustering) are optimized in totally different ways as summarized in Table I.
Original Space Discriminator (classify $x$ in the original space)
GAN [11]
Is $x$ from the ground truth or randomly generated?
@@ I-A Related works on adversarial net based X
$\left\{\begin{array}[]{c}if\;x\sim p(data),\;T=1\\ else\;x\sim p(x\mid z),\;T=0\end{array}\right.$
Latent Space Discriminator (classify sample $z$ in latent space)
AAE [12]
Is $z$ from a Gaussian prior?
$\left\{\begin{array}[]{c}if\;z\sim\mu+\sigma\cdot\mathcal{N}(0,1),\;T=1\\ else\;z\sim q(z\mid x),\;T=0\end{array}\right.$
DASC [13]
Does $z$ falls along the assigned subspace?
@@
$\left\{\begin{array}[]{c}if\;z\sim p(\varsigma^{*}),\;T=1\\ else\;z\sim p(\hat{\varsigma}),\;T=0\end{array}\right.$
DCAN (Ours)
Does $z$ belong to the assigned cluster?
$\left\{\begin{array}[]{c}if\;z\sim p(z\mid\theta^{*}),\;T=1\\ else\;z\sim q(z\mid x),\;T=0\end{array}\right.$
Table I: Different strategies of using adversarial net
II Background
II-A Kmeans based deep clustering
The simplest form of deep clustering uses the Euclidean distance loss in the Autoencoder based clustering (ABC) [7] approach. In ABC, the deep clustering objective optimizes the neural network to be affected by Kmeans clustering or $\mathcal{L}_{1}$ loss function in eqn (1). On the contrary, traditional reconstruction loss or $\mathcal{L}_{2}$ does not utilize any information from Kmeans.
$\begin{array}[]{c}\underset{w,b}{min}\;\left\{\mathcal{L}_{1}+\lambda\cdot\mathcal{L}_{2}\right\}\\ =\underset{w,b}{max}\;-\frac{1}{2}\left(T-y\right)^{2}-\lambda\left\{-\frac{1}{2}\left(h_{z}-\eta^{*}\right)^{2}\right\}\end{array}$ (1)
II-B KLD based deep clustering
Training a deep clustering algorithm such as ABC in eqn (1) requires a closed-form equation for the clustering loss function. In higher dimensional latent space, the Euclidean function will become less robust due to the curse of dimensionality. The probabilistic variant of ABC is based on the “KLD between two Gaussians” in [4, 9]. When the probabilistic variant is no longer a “KLD between two Gaussians”, the main problem is that:
1) the loss function can become non-trival to differentiate for backprogation.
Instead, we propose a new way to train deep clustering without facing such constrains i.e. we discard the use of such probabilistic function in deep clustering. Instead, we approximate deep clustering using adversarial net in two simple steps:
A) From ABC to KLD to JSD: We show that under certain condition (i.e. unit variance assumption), ABC is identical to KLD and thus related to JSD. Both KLD and JSD are also under the family of the $f$-divergence.
B) Adversarial net as JSD: We approximate JSD using adversarial net. This is possible since adversarial net approaches JSD when the training becomes optimal [11].
Our main goal is to find a relationship between adversarial net and deep clustering. First, we must first identify the relationship between KLD and ABC as seen in [4]. KLD measures the probabilistic distance between two distributions. The distributions of the latent space and deep clustering space [^1] We refer to deep clustering space as $p(z\mid\theta)$ and the latent space as $q(z\mid x)$. We refer to deep clustering space as the latent space partitioned using GMM. are defined as $q(z\mid x)$ and $p(z\mid\theta)$ respectively. The latent space is defined using the re-parameterization trick in [10] i.e. $z=\mu+\varrho\cdot\epsilon$. The deep clustering space is defined as $p(z\mid\theta)$, whereby $z=\left\{z^{(n)}\right\}_{n=1}^{N}\in\mathbb{R}^{Z}$ and $\theta=\{\eta,\tau,\zeta\}$, which are the mean $\eta$, precision $\tau$ and assignment parameters $\boldsymbol{\zeta}$ of GMM for $K$ number of clusters and $N$ number of samples. Also, $\eta=\left\{\eta_{k}\right\}_{k=1}^{K}\in\mathbb{R}^{Z},\tau=\left\{\tau_{k}\right\}_{k=1}^{K}\in\mathbb{R}^{Z},\zeta_{k}^{(n)}\in\{0,1\}$ and $\zeta^{(n)}$ is a $1-of-K$ vector. Thus, KLD based clustering loss in [4] as follows:
@@ II-B KLD based deep clustering
$\begin{array}[]{c}D_{KL}\left(\;p(z\mid\theta)\parallel q(z\mid x)\;\right),\\ p(z\mid\theta)=\prod_{k=1}^{K}\mathcal{N}(z\mid\eta_{k},\left(\tau_{k}\right)^{-1})^{\zeta_{k}},\\ q(z\mid x)=\mu+\sigma\cdot\mathcal{N}(0,1)\end{array}$ (2)
The problem is how do we define KLD in terms of two Gaussians, instead of a Gaussian and a GMM in eqn (4), which is intractable. To overcome this, we re-express the GMM term in $p(z\mid\theta)$ as follows
[…]
We define $p(z\mid\theta^{*})$ as the $k^{th}$ Gaussian component of Kmeans or GMM that generates sample $z$ in the latent space. i.e. we use $\zeta$ to compute $\eta^{*}$ and $\tau^{*}$ (refer to GMM in Appendix). Thus, the KLD of GMM and latent space in eqn (2) can now become a KLD between two Gaussian distributions in eqn (4). More importantly, a closed formed equation is available for the latter.
$\begin{array}[]{c}D_{KL}\left(\>\mathcal{N}(z\mid\eta^{*},\left(\tau^{*}\right){}^{-1})\parallel\mathcal{N}(z\mid\mu,\varrho^{2})\>\right)\end{array}$ (4)
The relationship between KLD and ABC can further be explained: If we discard the second order terms i.e. $\varrho^{2}$ and $\tau$ in eqn (4), the KLD loss reverts back to the original ABC loss [7, 14] in eqn (1).
III Proposed JSD based deep clustering
The problem with using JSD for deep clustering is that there is rarely a closed form solution available for a “JSD between two Gaussian distributions”. GAN [11] overcome this by employing an adversarial training procedure that approximates $D_{JS}\left(p_{data}\parallel p_{g}\right)$ at optimality, where $p_{data}$ is the real data distribution and $p_{g}$ is the generated data distribution. Unlike the discriminator of GAN in the original space, the discriminator of DCAN works in the latent space. Despite that, we can reformulate GAN into DCAN as follows:
$\begin{array}[]{c}D_{JS}\left(p(z\mid\theta^{*})\parallel q(z\mid x)\right)\\ =\underset{z\sim p(z\mid\theta^{*})}{E}\left[\ln D(z)\right]+\underset{x\sim p_{data}}{E}\left[\ln\left(1-D\left\{G(x)\right\}\right)\right]\\ =\underset{z\sim p(z\mid\theta^{*})}{E}\left[\ln D(z)\right]+\underset{z\sim q(z\mid x)}{E}\left[\ln\left(1-D\left\{z\right\}\right)\right]\end{array}$ (5)
Whereby $E_{z\sim p(z\mid\theta^{*})}\left[\cdot\right]$ refers to the expectation function where the sample is distributed from deep clustering space and $E_{z\sim q(z\mid x)}\left[\cdot\right]$ refers to the expectation function where the sample is distributed from the latent space.
IV Training the Proposed Method
We can visualize $D_{JS}\left(\>p(z\mid\theta^{*})\parallel q(z\mid x)\>\right)$ using Fig. 1. The discriminator tries to distinguish samples from deep clustering space, $z\sim p(z\mid\theta^{*})$ versus samples generated from the encoder output, $z\sim q(z\mid x)$. We train DCAN by using the two backward paths shown by the colored dashed lines in Fig 1. We attach a pseudocode of DCAN in Algorithm 1.
Neural Network B (blue dash line) refers to the backward path from cross-entropy loss to latent space input. The discriminator weight is updated by this path.
Neural Network R (red dash line) refers to the backward path from cross-entropy loss to image space input. We use this path to update both discriminator and encoder weight.
@@ IV Training the Proposed Method
Figure 1: Proposed deep clustering loss using adversarial net. It uses the discriminator and encoder architecture inspired by GAN.
IV-A Training the discriminator
For discussion purpose, we only use a one hidden layer MLP for the encoder and the discriminator. Extending it to two or more hidden layers is straightforward. We use a MLP structure $x_{i}-h_{j}-z_{\varsigma}$ for our encoder and $z_{\varsigma}-h_{f}-y_{d}$ for our discriminator. The hidden layers are defined as $h_{f}$, $h_{j}$ and $z_{\varsigma}$ (encoder output) respectively. We train the discriminator weight $\left\{w_{fd},w_{\varsigma f}\right\}$ by feeding it with $1...N$ positive and negative samples. For the sake of brevity, we refer to each $n^{th}$ sample of $x$ as $x^{\left(n\right)}$ with dimension $I$. In the forward pass in Fig 1, samples go to the discriminator through path B and R. Samples generated from deep clustering space1 (path B) are positive i.e. $T=1$ and samples generated from latent space (path R) are negative i.e. $T=0$. Cross-entropy loss is defined for $L_{2}$ and $L_{1}$ using the following:
$\begin{array}[]{c}D=y^{T}(1-y)^{1-T}\\ L_{1}=\ln D(z)\\ L_{2}=\ln\left[1-D(G(x))\right]\end{array}$ (6)
The activation function for each layer is represented as $\sigma\left(\cdot\right)$. We define the neural network output and target label as $y=\left\{0,1\right\}$ and $T=\left\{0,1\right\}$ respectively. Both $y$ and $T$ have the same dimension $d$.
Neural Network R: We perform weight update using eqn (7) via stochastic gradient ascent (SGA). As we seek sampling from $p(z\mid\theta^{*})$ instead of $q(z\mid x)$, we set $T=0$ for $D$.

### v3

@@ Abstract
Deep clustering is a recent deep learning technique which combines deep learning with traditional unsupervised clustering. At the heart of deep clustering is a loss function which penalizes samples for being an outlier from their ground truth cluster centers in the latent space. The probabilistic variant of deep clustering reformulates the loss using $KL$ divergence. Often, the main constraint of deep clustering is the necessity of a closed form loss function to make backpropagation tractable. Inspired by deep clustering and adversarial net, we reformulate deep clustering as an adversarial net over traditional closed form $KL$ divergence. Training deep clustering becomes a task of minimizing the encoder and maximizing the discriminator. At optimality, this method theoretically approaches the $JS$ divergence between the distribution assumption of the encoder and the discriminator. We demonstrated the performance of our proposed method on several well cited datasets such as MNIST, REUTERS and CIFAR10, achieving on-par or better performance with some of the state-of-the-art deep clustering methods.
1 Introduction
@@
Deep neural network such as the autoencoder is applied to many signal processing domains such as speech and audio processing [1, 2], image clustering [3, 4] and medical data processing [5, 6]. While the latent space of autoencoder is well suited for dimension reduction through reconstruction loss [7], the latter is not optimized for clustering/ classification since class labels cannot be used in the reconstruction loss [8]. A recent autoencoder technique known as the Euclidean distance based clustering (ABC) [7], utilize class label information in neural network by introducing a loss function known as the deep clustering loss. The goal is to minimizes the Euclidean distance in the latent space between samples and partitioning learnt by clustering algorithms e.g. Kmeans or Gaussian mixture model (GMM). In other words, the latent space of an autoencoder learnt using reconstruction loss will look different when we further perform K-means on the latent space. A recent probabilistic approach to ABC uses KL divergence (KLD) [4, 9] and assumes that both the variational autoencoder (VAE) [10] latent space and clustering approach are Gaussian distributed. We can compare the similarities and differences between deep clustering and VAE to better understand the former. The neural network of VAE and deep clustering are optimized by backpropagating samples in the latent space to update the encoder weights. Also, both deep clustering and VAE share a similar goal of modeling each input image as a sample draw of the Gaussian distribution in the latent space. However, their Gaussian distributions are different. VAE enforce samples in the latent space to be closely distributed by a Gaussian distribution with continuous mean $\mu$ and variance $\sigma$ i.e. $z\sim\mu+\sigma\cdot\mathcal{N}(0,1)$. Whereas the latent space of deep clustering approaches a Gaussian mixture model (GMM) and each Gaussian has a set of fixed parameters i.e. $z\mathcal{\sim N}(\eta_{k},\tau_{k})$, $\forall k\in K$. The goal of VAE is to enforce all class samples in the latent space to be distributed by continuous mean and variance. The goal of deep clustering is to enforce class samples in the latent space to be distributed by a discrete set of mean and variance i.e. $N(\eta^{*},\tau^{*})$. Thus, only deep clustering is associated with clustering, making it more suitable for unsupervised classification. Deep clustering is not without problem. Mainly, when backpropagating from a neural network, we require a closed form loss function for tractability [11]. Most deep learning methods including VAE restrict to the assumption of a KLD between two Gaussians to obtain a closed form loss function for backpropagating. When generalizing to $f$-divergence between two Gaussians, closed form solution does not exist. A well known workaround to this problem is to approximate the JS divergence (JSD) between two Gaussians using adversarial net. In fact, adversarial net is quite versatile as shown in Table 1. For different applications, the discriminator can be seen as performing a specific task different from the discriminators in other adversarial net. We distinguish the proposed work from other “adversarial net based X” in Table I where X can be any discriminator task. In GAN, X refers to image generation. Other instances of “adversarial net based X” includes AAE, DASC and etc. In our approach, “Deep clustering using adversarial net based clustering loss” (DCAN), X refers to clustering. More specifically, in DCAN the JSD between latent space and clustering is approximated by adversarial net. More details are discussed in the next section.
Adversarial net based X Discriminator
@@
GAN [11] $\left\{\begin{array}[]{c}if\;x\sim p(data),\;T=1\\ else,\;T=0\end{array}\right.$ $x$ is from the dataset?
@@
DASC [12] $\left\{\begin{array}[]{c}if\;z\sim p(\varsigma^{*}),\;T=1\\ else,\;T=0\end{array}\right.$ $z$ is from an assigned subspace?
@@
AAE [13] $\left\{\begin{array}[]{c}if\;z\sim\mu+\sigma\cdot\mathcal{N}(0,1),\;T=1\\ else,\;T=0\end{array}\right.$ $z$ is from a Gaussian prior?
DCAN (proposed) $\left\{\begin{array}[]{c}if\;z\sim p(z\mid\theta^{*}),\;T=1\\ else,\;T=0\end{array}\right.$ $z$ is from an assigned cluster?
Table 1: Different strategies of using adversarial net based X
1.1 Adversarial net based X
The optimization of adversarial net [11] in general centers around minimizing the encoder/generator and maximizing the discriminator. However, as mentioned “adversarial net based X” are versatile and we highlight their similarities and differences as follows: In GAN, the generator is responsible for generating an image when given a sample from the latent space. The discriminator of GAN checks whether the generated image belongs to the distribution of the original dataset. Similarly, in DCAN we minimize the encoder and maximize the discriminator but for different purpose. In DCAN, the encoder receives an image and outputs a sample in the latent space. GAN trains the generator to be sensitive to the variance of the image class i.e. small displacement of the sample in the latent space can result in huge image variance. DCAN trains the encoder to be insensitive to the class variance i.e. different images from the same class populate tightly close to the cluster center in the latent space. Both DCAN and AAE [13] utilize the adversarial net in totally different ways despite having similar architecture (encoder acting as generator). The goal of AAE is identical to VAE i.e. the AAE encoder models a sample in the latent space as a sum and product of $\mathcal{N}(0,1)$. On the other hand, the goal of DCAN is not identical to VAE. DCAN models a sample in the latent space, as the sample draw from a cluster (e.g. Kmeans or Gaussian mixture model), specifically the sample draw of a $1-of-K$ cluster. The goal of DASC [12] is subspace clustering, and it is different from deep clustering. Like principal component analysis, DASC perform unsupervised learning by decomposing the dataset into different individual eigenvectors, each eigenvector as orthogonal to each other as possible. Samples from each classes should ideally reside on their respective $K^{th}$ subspaces, just as each sample should ideally fall within their respective $K^{th}$ clusters in deep clustering. In DASC, the latent space of DASC is trained to behave like subspace clustering. Thus the adversarial net of all four methods, GAN (models the dataset), AAE (models VAE loss), DASC (models subspace clustering) and DCAN (models deep clustering) are optimized in totally different ways as summarized in Table I.
2 Proposed method
@@
Figure 1: Proposed deep clustering loss using adversarial net. It uses the discriminator and encoder architecture inspired by AAE.
One motivation for using a JSD approach is that it does not suffer from asymmetry unlike KLD [14, 11]. However, there is no closed form solution available for the JSD between two data distributions. GAN [11] overcome the need for a closed form solution by employing an adversarial training procedure. GAN approximates JSD or $D_{JS}\left(p_{data}\parallel p_{g}\right)=E_{x\sim p_{data}}\left[\log D_{G}(x)\right]+E_{x\sim p_{g}}\left[\log\left(1-D_{G}\left\{x\right\}\right)\right]$ at optimality, where $p_{data}$ is the real data distribution and $p_{g}$ is the generated data distribution. In the JSD between two Gaussian distributions for deep clustering, we seek $D_{JS}\left(\>p(z\mid\theta^{*})\parallel q(z\mid x)\>\right)=E_{z\sim p(z\mid\theta^{*})}\left[\ln D(z)\right]+E_{z\sim q(z\mid x)}\left[\ln\left(1-D\left\{z\right\}\right)\right]$, whereby $z\sim p(z\mid\theta^{*})$ refers to the sample distributed from clustering and $z\sim q(z\mid x)$ refers to the sample distributed from the encoder. Unlike the discriminator of GAN in the original space, the discriminator of DCAN works in the latent space. We can visualize $D_{JS}\left(\>p(z\mid\theta^{*})\parallel q(z\mid x)\>\right)$ using Fig. 1. The discriminator tries to distinguish samples from clustering, $z\sim p(z\mid\theta^{*})$ versus samples generated from the encoder, $z\sim q(z\mid x)$.
2.1 Training DCAN
We train DCAN by feeding it with $1...N$ positive and negative samples. We refer to each $n^{th}$ sample of $x$ as $x^{\left(n\right)}$ with dimension $I$. In the forward pass in Fig 1, samples enter the discriminator through the encoder and clustering. Samples generated from clustering1 are positive i.e. $T=1$ and samples generated from encoder are negative i.e. $T=0$. We define the neural network output and target label as $y=\left\{0,1\right\}$ and $T=\left\{0,1\right\}$ respectively. Both $y$ and $T$ have the same dimension $d$. We update the discriminator weight using DCAN loss in eqn (1). Since the samples in the latent space displaces each time the encoder weight changes, clustering parameters $\theta_{k}=\{\eta_{k},\tau_{k},\pi^{(n)}\forall k\}$ has to be re-computed each time the discriminator weight is updated. The encoder $q(z\mid x)$ maps $x$ to $z$ in the latent space. However, we desire to minimize between $q(z\mid x)$ and $p(z\mid\theta^{*}).$ Meaning that we desire to produce a sample $z$ from $q(z\mid x)$ that is as close as possible to the sampling from $p(z\mid\theta^{*})$. Similarly, DCAN loss in eqn (1) is used to update the encoder weight.